# NoLiMa: Long-Context Evaluation Beyond Literal Matching

### v2

@@ 4.3 Evaluation Setup & Metric
Figure 2: The full sweep plots (a & b) illustrate performance across the entire context window. The plots for the last 2K tokens (c & d) depict performance when needle placements are aligned within the final 2K tokens for various context lengths. The color shading of each plot line represents the tested context length. To minimize noise and highlight trends more clearly, we increased the number of placements from 26 to 51 and applied a moving average with a window size of 12.
4.4 Results
Table 3 presents the performance results of all NoLiMa tests on the selected models. Most models achieve high base scores, indicating that the designed needle set is relatively simple to answer in shorter contexts. Even models with base scores exceeding 90.0% exhibit a significantly shorter effective length than their claimed lengths, generally limited to $\leq$2K tokens, with GPT-4o being an exception. While GPT-4o demonstrates strong overall performance, it fails to generalize effectively beyond 8K tokens. Out of the 12 models, 10 exhibit performance at 32K lengths that is half or less of their base scores. For comparison, in other benchmarks with similar settings, such as BABILong (QA1) (Kuratov et al., 2024) and RULER (Hsieh et al., 2024), Llama 3.1 70B achieves effective lengths of 16K [^4] In BABILong, the effective length is also based on 85% of the 0K base performance threshold and 32K, respectively. However, in NoLiMa, Llama 3.1 70B has an effective length of only 2K and shows a significant drop in performance at 32K lengths (42.7% vs. 94.3% base score). Models such as Claude 3.5 Sonnet, Gemini 1.5 Flash, GPT-4o mini, and Llama 3.1 8B may have weaker base scores, but their effective lengths are calculated relative to these scores. This reveals an interesting observation: a model like Claude 3.5 Sonnet, despite having a lower base score, may underperform in shorter contexts but demonstrate better length generalization than models with higher base scores, such as Llama 3.1 70B and Llama 3.3 70B. In fact, Sonnet even achieves higher raw scores in 4K-token experiments compared to some higher-base-score models.
4.4.1 Latent hops & Inversion
As discussed in Section 3, our needle set also includes examples requiring two-hop associative linking from the question keyword to the needle keyword. To evaluate the impact on length generalization, Figure 3(a) presents the normalized performance of two top-performing models on one-hop and two-hop tasks. It is evident that, for the same context lengths, questions involving two-hop latent reasoning steps are more challenging than those requiring one-hop reasoning. Notably, the performance gap between one-hop and two-hop tasks widens with increasing context lengths. GPT-4o demonstrates impressive generalization performance, handling both types of examples effectively even at context lengths up to 4K.
Each group of needles includes both a default and an inverted template and Figure 3(b) shows that inverted examples are more challenging to answer. We argue this difficulty arises from the model’s causal attention mechanism, particularly in longer contexts where attention signals weaken. In the default template, the question or particularly $W_{q}$, can link directly to $W_{n}$, which could contain information about the character’s name since the name appears earlier in the sequence. This allows the model to backtrace effectively from $W_{q}$ through $W_{n}$ to the character. In the inverted template, $W_{q}$ may still attend to $W_{n}$, but since the fact is incomplete (the character hasn’t been stated yet), the model cannot use that attention to resolve the question. Instead, it must rely on weaker signals encoded in the character’s name to establish the link, which becomes harder with longer contexts due to diminishing attention strength. While these findings shed light on the challenge, deeper mechanistic analysis is beyond the scope of this paper and requires further study.
(a) One-hop vs. Two-hop
(b) Default vs. Inverted
Figure 3: Impact of (a) number of hops and (b) inversion on normalized performance across GPT-4o and Llama 3.3 70B models. The red dotted line indicates the 0.85 effective threshold.
@@ 4.4.2 Needle Placement Depth Analysis
A common evaluation across NIAH-based benchmarks (Kamradt, 2023) examines the impact of needle placement within the context window. In Figure 2(a), we observe a ”lost-in-the-middle” effect (Liu et al., 2024) in 32K, where model performance dips when the needle appears in the middle of longer contexts.
Additionally, Figure 2(b) reveals a key phenomenon: longer contexts in more complex (two-hop) examples dampens the performance distribution over the full sweep depending on their length. In vanilla multi-document or NIAH-based benchmarks (Kamradt, 2023; Liu et al., 2024), models perform consistently well when the needle (or gold document) appears at the very beginning or end of the context window, with minimal impact from context length. However, in NoLiMa, as task complexity increases in two-hop scenarios, larger context sizes shift the entire trendline downward toward zero, with performance declining even at the edges of the context window.
To further investigate this issue, we devise an alternative setup that focuses on analyzing the last 2K tokens instead of sweeping across the full context. Therefore, we align the placement positions in the last 2K tokens for all context lengths (see Figure 4). This makes that for a certain token depth the only changing factor in each plotline would be the context length, which in turn means that the model has more tokens that can be attended to.
Based on the final 2K results in Figure 2(c), the one-hop setup confirms our earlier observations from the full-sweep plots. The “lost-in-the-middle” phenomenon—where performance dips toward the center of the context—primarily appears in simpler tasks. Each plotline drops as it moves toward the center, reflecting its dependence on placement position and the way the model encodes positional information. In contrast, the two-hop scenario appears to be influenced more by attention limitations than by position encoding alone. Figure 2(d) reveals that, rather than depth exacerbating performance drops, the plot lines remain relatively stable over the last 2K positions. However, context length significantly reduces the overall performance trends observed in this range. Llama 3.x models, like many other recent language models, features rotary position embeddings (RoPE) which is a relative PE (Su et al., 2024). For each token depth in Figure 2(d), as the relative distance between question and fact remains the same regardless of context length, position encoding does not explain the performance drop. Instead, the main limiting factor is the increased context length: as the number of tokens grows, the attention mechanism struggles to process information effectively. In the absence of strong surface-level cues (e.g., literal matches), locating relevant facts becomes challenging for the model, regardless of their position within long contexts.
@@ 4.4.4 Ablation Study: Literal match effect
Distracting Literal Matches.
While literal matches could serve as cues if they are part of the relevant fact, they can also act as distractors if they are irrelevant to the answer. In Section 2, we noted that some related benchmarks include similar documents in the context as distractors to test the model’s ability to discern the correct answer from irrelevant ones. This setup creates matches between the query and both relevant and irrelevant documents or facts. In contrast, NoLiMa allows us to explore a different scenario: when the context contains distracting words overlapping with the question, while the relevant fact has minimal overlap with the query. We insert a distractor sentence into the haystack (details in Appendix D) containing $W_{q}$ but entirely irrelevant to both the needle and the question’s intent. This setup poses a significant challenge, requiring the model to disregard irrelevant literal overlaps while identifying a relevant fact with no meaningful overlap with the query. As shown in Figure 5, such distractors have a substantial impact on degrading length generalization. GPT-4o now demonstrates an effective length of just 1K, while Llama 3.3 70B performs even worse. While adding distractors slightly lowers base scores (GPT-4o: 93.8, Llama 3.3 70B: 84.4), the normalized plots still clearly illustrate a performance drop at longer lengths. These results highlight the challenge of resolving queries in contexts where irrelevant overlaps mislead the model, and the relevant fact shares no overlap with the question.
@@ 5 Conclusion
NoLiMa provides a challenging benchmark for evaluating the reasoning capabilities of large language models in long-context settings. By removing literal overlaps between questions and relevant information, the benchmark tests models’ ability to infer and link information within extensive irrelevant content. Our findings show that even state-of-the-art models struggle, especially as context length increases, revealing serious limits in their attention mechanism. While causal attention should theoretically access all previous tokens, models often rely on surface-level cues in longer contexts. This vulnerability becomes more pronounced when the context contains literal matches that fail to connect with the truly relevant fact, causing models to overlook the correct information and focus instead on superficial signals. We believe our findings with NoLiMa are likely to extend to downstream applications. For instance, in search engines or RAG systems, a relevant document containing the correct answer may have a lexical gap with the query. So, even if such a document is retrieved alongside others that likely have higher lexical similarity, language models may struggle to extract the correct answer, as they can become distracted by the lexical overlaps in those other documents. This work highlights the need for benchmarks that go beyond surface-level retrieval to assess deeper reasoning. NoLiMa sets a new standard for evaluating long-context comprehension and emphasizes the importance of developing approaches capable of handling complex reasoning in long contexts.
@@ Acknowledgments
We thank Abdullatif Köksal, Leonie Weissweiler, and Amir Hossein Kargaran for their valuable feedback and support, particularly in the early stages of this project.

### v3

@@ 4.4.1 Latent hops & Inversion
Figure 3: The full sweep plots (a & b) illustrate performance across the entire context window, where 0% corresponds to the beginning of the haystack and 100% to the end. The plots for the last 2K tokens (c & d) depict performance when needle placements are aligned within that range for various context lengths; 0 marks the end of the context, and larger values indicate positions farther from the end (up to 2K tokens inward). The color shading of each plot line represents the tested context length. To minimize noise and highlight trends more clearly, we increased the number of placements from 26 to 51 and applied a moving average with a window size of 12.666All figures use Llama 3.3 70B. Plots without smoothing applied are available in Appendix G.
@@ 4.4.2 Needle Placement Depth Analysis
A common evaluation across NIAH-based benchmarks (Kamradt, 2023) examines the impact of needle placement within the context window. In Figure 3(a), we observe a “lost-in-the-middle” effect (Liu et al., 2024) in 32K, where model performance dips when the needle appears in the middle of longer contexts.
Additionally, Figure 3(b) reveals a key phenomenon: longer contexts in more complex (two-hop) examples dampen the performance distribution over the full sweep depending on their length. In vanilla multi-document or NIAH-based benchmarks (Kamradt, 2023; Liu et al., 2024), models perform consistently well when the needle (or gold document) appears at the very beginning or end of the context window, with minimal impact from context length. However, in NoLiMa, as task complexity increases in two-hop scenarios, larger context sizes shift the entire trendline downward toward zero, with performance declining even at the edges of the context window.
To further investigate this issue, we devise an alternative setup that focuses on analyzing the last 2K tokens instead of sweeping across the full context. Therefore, we align the placement positions in the last 2K tokens for all context lengths (see Figure 4). This ensures that for a certain token depth the only changing factor in each plotline is the context length, which in turn means that the model has more tokens that it needs to attend to.
Based on the final 2K results in Figure 3(c), the one-hop setup confirms our earlier observations from the full-sweep plots. The “lost-in-the-middle” phenomenon—where performance dips toward the center of the context—primarily appears in simpler tasks. Each plotline drops as it moves toward the center, reflecting its dependence on placement position and the way the model encodes positional information. In contrast, the two-hop scenario appears to be influenced more by attention limitations than by position encoding alone. Figure 3(d) reveals that, rather than depth exacerbating performance drops, the plot lines remain relatively stable over the last 2K positions. However, context length significantly reduces the overall performance trends observed in this range. Llama 3.x models, like many other recent language models, feature rotary position embeddings (RoPE), a relative PE (Su et al., 2024). For each token depth in Figure 3(d), as the relative distance between question and fact remains the same regardless of context length, position encoding does not explain the performance drop. Instead, the main limiting factor is the increased context length: as the number of tokens grows, the attention mechanism struggles to process information effectively. In the absence of strong surface-level cues (e.g., literal matches), locating relevant facts becomes challenging for the model, regardless of their position within long contexts.
4.4.3 CoT Prompting
@@ 4.4.4 Ablation Study: Literal match effect
Distracting Literal Matches.
While literal matches serve as cues if they are part of the relevant fact, they can also act as distractors if they are irrelevant to the answer. In Section 2, we noted that some related benchmarks include similar documents in the context as distractors to test the model’s ability to discern the correct answer from irrelevant ones. This setup creates matches between the query and both relevant and irrelevant documents or facts. In contrast, NoLiMa allows us to explore a different scenario: when the context contains distracting words overlapping with the question, while the relevant fact has minimal overlap with the query. We insert a distractor sentence into the haystack (details in Appendix D) containing $W_{q}$ but entirely irrelevant to both the needle and the question’s intent. This setup poses a significant challenge, requiring the model to disregard irrelevant literal overlaps while identifying a relevant fact with no meaningful overlap with the query. As shown in Figure 5, such distractors have a substantial impact on degrading length generalization. GPT-4o now demonstrates an effective length of just 1K, while Llama 3.3 70B performs even worse. While adding distractors slightly lowers base scores (GPT-4o: 93.8, Llama 3.3 70B: 84.4), the normalized plots still clearly illustrate a performance drop at longer lengths. These results highlight the challenge of resolving queries in contexts where irrelevant overlaps mislead the model, and the relevant fact shares no overlap with the question.
@@ 5 Conclusion
NoLiMa provides a challenging benchmark for evaluating the reasoning capabilities of large language models in long-context settings. By removing literal overlaps between questions and relevant information, the benchmark tests models’ ability to infer and link information within extensive irrelevant content. Our findings show that even state-of-the-art models struggle, especially as context length increases, revealing serious limits in their attention mechanism. While causal attention should theoretically access all previous tokens, models often rely on surface-level cues in longer contexts. This vulnerability becomes more pronounced when the context contains literal matches that fail to connect with the truly relevant fact, causing models to overlook the correct information and focus instead on superficial signals. We believe our findings with NoLiMa are likely to extend to downstream applications. For instance, in search engines or RAG systems, a relevant document containing the correct answer may have a lexical gap with the query. So, even if such a document is retrieved alongside others that likely have higher lexical similarity, language models may struggle to extract the correct answer, as they can become distracted by the lexical overlap with these other documents. This work highlights the need for benchmarks that go beyond surface-level retrieval to assess deeper reasoning. NoLiMa sets a new standard for evaluating long-context comprehension and emphasizes the importance of developing approaches capable of handling complex reasoning in long contexts.
@@ Acknowledgments
We thank Abdullatif Köksal, Leonie Weissweiler, and Amir Hossein Kargaran for their valuable feedback and support, particularly in the early stages of this project. We also appreciate the insights from our peers, and we are grateful to the anonymous reviewers for their constructive comments.
@@ Appendix B Models
Llama 4 Maverick 1M Yes meta-llama/Llama-4-Maverick-17B-128E-Instruct
Llama 4 Scout 10M Yes meta-llama/Llama-4-Scout-17B-16E-Instruct
@@
Gemini 2.0 Flash 1M No gemini-2.0-flash
Gemini 2.5 Flash 1M No gemini-2.5-flash-preview-05-20
Gemma 3 27B 128K Yes google/gemma-3-27b-it
Gemma 3 12B 128K Yes google/gemma-3-12b-it
Gemma 3 4B 128K Yes google/gemma-3-4b-it
@@ Appendix D Distractor Design
Models Claimed Effective Base Score 1K 2K 4K 8K 16K 32K 64K 128K777127,500-token haystack used based on model token limit constraints.
Length Length ($\times$0.85: Thr.)
GPT-4.1 1M 16K 97.0 (82.5) 95.6 95.2 91.7 87.5 84.9 79.8 69.7 64.7
GPT-4o 128K 8K 99.3 (84.4) 98.1 98.0 95.7 89.2 81.6 69.7 62.4 56.0
Gemini 2.5 Flash 1M 2K 94.4 (80.2) 90.1 86.1 79.4 68.2 57.9 48.4
Gemini 2.0 Flash 1M 4K 89.4 (76.0) 87.7 87.5 77.9 64.7 48.2 41.0 33.0 16.4
Llama 4 Maverick 1M 2K 90.1 (76.6) 81.6 78.3 68.8 49.0 34.3 24.5
Gemma 3 27B 128K $<$1K 88.6 (75.3) 73.3 65.6 48.1 32.7 20.2 9.5
Gemma 3 12B 128K 1K 87.4 (74.3) 74.7 61.8 39.9 27.4 16.8 7.3
Llama 4 Scout 10M 1K 81.7 (69.4) 72.3 61.8 50.8 35.5 26.9 21.6
GPT-4.1 Mini 1M $<$1K 80.9 (68.8) 66.7 62.8 58.7 51.9 46.2 38.8
GPT-4.1 Nano 1M $<$1K 80.7 (68.6) 60.8 48.2 36.7 28.8 19.5 9.4
Gemma 3 4B 128K $<$1K 73.6 (62.6) 50.3 35.3 16.4 7.5 2.3 0.9
Table 10: NoLiMa benchmark results for GPT-4o, Gemini 2.0 Flash, and additional recent models. For models with stronger performance at 32K, we extend evaluation to 64K and 128K using the same setup. Scores above the effective threshold are underlined; scores below 50% of the base score are shaded in red.
Appendix E Results Beyond 32K & Recent LLMs
Based on the dataset configuration outlined in Section 4.1, we construct haystacks by randomly concatenating snippets extracted from the filtered books dataset. This setup enables evaluation across scalable context lengths, including 64K and 128K tokens, and can be extended further as model limits allow.
Two practical adjustments were made in this evaluation setup, without altering the core methodology: (1) Reduced Placement Count: Due to cost limitations, we reduce the number of needle placements from the default 26 to 11 placements per context length. This change minimizes API usage while still preserving meaningful coverage across the haystack. (2) Token Limit Constraints: GPT-4o has a strict 128,000 token limit, including both input and output tokens. To accommodate the task prompt and model response, we limit the haystack length to 127,500 tokens for this model. To ensure a fair comparison, we use the same haystack length for Gemini 2.0 Flash.
The results of this evaluation are presented in Table 10, focusing on GPT-4o and Gemini 2.0 Flash at 64K and 128K context lengths. While GPT-4o does not fully meet its claimed performance at the maximum context length, it nonetheless maintains over 50% of its base score at 128K, indicating relatively strong performance at that scale. In contrast, Gemini 2.0 Flash shows a sharper decline, dropping to 16.4% at 128K.
To cover recent model releases, Table 10 also includes entries for newer models such as the GPT-4.1 series and Gemini 2.5 Flash [^8] Without reasoning (thinking budget = 0). (Team, 2025; Gemini Team et al., 2024). For those that showed stronger performance at 32K, we also extend the evaluation to 128K using the same setup. GPT-4.1 shows clear improvements over prior models; however, its effective context length remains around 16K–well below the claimed 1M–and its performance drops below 65% on 128K context lengths.
Appendix F One- & Two-hop Results
Tables 11 and 12 present detailed results on the one-hop and two-hop subsets of the NoLiMa benchmark, evaluated across the selected models. The tables follow the same format and thresholding criteria as Table 3, reporting both the claimed and effective context lengths. Note that the base scores–and consequently the thresholds–are computed separately for each subset. Although both one-hop and two-hop subsets show strong base scores, two-hop tasks generally yield shorter effective context lengths. This suggests that while models can perform well on complex reasoning in short contexts, their performance degrades more rapidly as context length increases, indicating reduced length generalization under greater reasoning demands.
Models Claimed Effective Base Score 1K 2K 4K 8K 16K 32K
Length Length ($\times$0.85: Thr.)
GPT-4o 128K 16K 99.3 (84.4) 97.7 97.5 95.6 91.9 87.3 79.8
Llama 3.3 70B 128K 8K 97.7 (83.1) 97.1 93.6 90.3 84.1 73.2 56.2
Llama 3.1 405B 128K 4K 91.7 (78.0) 88.7 87.3 80.2 68.4 59.6 49.4
Llama 3.1 70B 128K 4K 96.4 (82.0) 95.2 89.8 82.7 76.9 66.3 56.9
Gemini 1.5 Pro 2M 4K 90.8 (77.2) 85.7 85.5 81.9 72.3 63.4 55.1
Jamba 1.5 Mini 256K 2K 93.1 (79.1) 80.0 80.2 77.9 71.5 62.9 56.0
Command R+ 128K 2K 91.6 (77.8) 79.4 78.4 75.6 52.2 26.9 10.4
Gemini 2.0 Flash 1M 4K 92.5 (78.6) 91.5 93.5 89.6 78.4 61.8 52.8
Mistral Large 2 128K 4K 83.3 (70.8) 82.5 86.1 80.3 62.5 44.1 27.5
Claude 3.5 Sonnet 200K 8K 90.5 (76.9) 89.9 91.6 89.5 78.2 61.1 45.2
Gemini 1.5 Flash 1M 1K 85.7 (72.9) 76.4 71.8 63.6 57.0 48.7 41.3
GPT-4o mini 128K 1K 88.4 (75.2) 81.0 73.6 57.6 45.4 30.2 20.0
Llama 3.1 8B 128K 1K 83.0 (70.5) 75.7 69.3 60.7 49.6 35.7 22.7
Table 11: NoLiMa benchmark results on one-hop examples. Base scores and effective lengths are computed using only the one-hop subset. Scores above the effective threshold are underlined, while scores that are below 50% of the base score are shaded in red.
Models Claimed Effective Base Score 1K 2K 4K 8K 16K 32K
Length Length ($\times$0.85: Thr.)
GPT-4o 128K 8K 99.3 (84.4) 98.7 98.7 95.8 85.9 74.6 57.4
Llama 3.3 70B 128K 1K 96.7 (82.2) 90.6 79.8 70.7 57.4 42.7 25.9
Llama 3.1 405B 128K 2K 95.3 (81.0) 89.4 82.0 67.4 49.8 34.6 23.8
Llama 3.1 70B 128K 1K 92.0 (78.2) 85.9 72.0 57.0 45.2 33.8 26.4
Gemini 1.5 Pro 2M 1K 94.9 (80.7) 87.1 79.4 67.4 53.6 45.7 39.6
Jamba 1.5 Mini 256K $<$1K 91.7 (77.9) 71.8 66.6 62.0 50.7 40.0 28.4
Command R+ 128K $<$1K 90.0 (76.5) 74.0 67.4 54.7 23.8 14.3 3.8
Gemini 2.0 Flash 1M 2K 85.7 (72.8) 83.1 80.0 63.6 47.8 31.4 26.4
Mistral Large 2 128K 2K 93.6 (79.5) 90.4 84.7 64.7 37.8 18.4 7.9
Claude 3.5 Sonnet 200K 2K 84.0 (71.4) 79.8 74.5 63.0 41.5 26.9 10.9
Gemini 1.5 Flash 1M $<$1K 83.5 (71.0) 59.1 49.0 35.6 28.8 19.2 12.9
GPT-4o mini 128K $<$1K 80.4 (68.3) 51.4 39.3 27.6 16.9 8.8 5.9
Llama 3.1 8B 128K $<$1K 68.9 (58.6) 53.4 36.0 23.6 10.0 6.5 3.8
Table 12: NoLiMa benchmark results on two-hop examples. Base scores and effective lengths are computed using only the two-hop subset. Scores above the effective threshold are underlined, while scores that are below 50% of the base score are shaded in red.
(a) Full Sweep (One-hop)
(b) Full Sweep (Two-hop)
(c) Last 2K (One-hop)
(d) Last 2K (Two-hop)
Figure 6: Unsmoothed needle placement depth plots corresponding to the smoothed results in Figure 6. These plots reflect raw performance values prior to applying the moving average.
Appendix G Raw Needle Placement Depth Plots
Figure 6 presents the needle placement depth plots corresponding to Figure 6, prior to the application of the moving average employed in the main figure.